# Quantum Entanglement Geometry on Severi–Brauer Schemes: Subsystem Reductions of Azumaya Algebras

### v1

@@ Abstract.
We formulate pure-state entanglement in families as a geometric obstruction. In standard quantum information, entanglement is defined relative to a chosen tensor-product factorization of a fixed Hilbert space. In contrast, for a twisted family of pure-state spaces, which can be described by Azumaya algebras $\mathcal{A}$ of degree $n$ on $X$ and their Severi–Brauer schemes
$SB(\mathcal{A})=P\times^{\operatorname{PGL}_{n}}\mathbb{P}^{n-1}\to X,$
such a subsystem choice may fail to globalize. We formalize this algebro-geometrically: fixing a factorization type $\mathbf{d}=(d_{1},\dots,d_{s})$ with $n=\prod_{i}d_{i}$, the existence of a global product-state locus of type $\mathbf{d}$ is equivalent to a reduction of the underlying $\operatorname{PGL}_{n}$–torsor $P\to X$ to the stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$. Thus, entanglement is the obstruction to the existence of a relative Segre subscheme inside $SB(\mathcal{A})$.
Writing $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ for the Segre variety, we call a reduction to $G_{\mathbf{d}}$ a $\mathbf{d}$–subsystem structure. Our first main result identifies the moduli of $\mathbf{d}$–subsystem structures with the quotient $P/G_{\mathbf{d}}$. Moreover, we realize naturally $P/G_{\mathbf{d}}$ as a locally closed subscheme of the relative Hilbert scheme,
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr)\ \subset\ \operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr),$
parametrizing relative closed subschemes fppf-locally isomorphic to $\Sigma_{\mathbf{d}}\times X$.
In the bipartite case $\mathbf{d}=(d_{A},d_{B})$, $G_{\mathbf{d}}$–invariant determinantal geometry globalizes the Schmidt-rank filtration, with base-change compatible resolutions and numerical invariants that are constant on fibers. Finally, we study Brauer-theoretic constraints on the existence of subsystem structures, including an $\mathrm{lcm}(\mathbf{d})$–torsion obstruction and examples showing that $\mathbf{d}$–reducibility depends on the underlying $\operatorname{PGL}_{n}$–torsor, not only on the Brauer class.
@@ 1. Introduction
which is precisely the locus of product states of type $\mathbf{d}=(d_{1},\dots,d_{s})$. In the bipartite case, the determinantal rank loci recover the Schmidt-rank stratification and provide a geometric entanglement filtration.
@@
which is fppf-locally a trivial $\mathbb{P}^{n-1}$–bundle. From the quantum-information viewpoint, $SB(\mathcal{A})\to X$ is a natural algebro–geometric model for a family of pure-state spaces whose projective trivializations may carry nontrivial twisting.
@@
    When does $SB(A)$ contain a relative Segre subscheme of type $\mathbf{d}$ (when does the associated $\mathrm{PGL}_{n}$-torsor admit a reduction to $G_{\mathbf{d}}$)? In terms of quantum information, when does $SB(\mathcal{A})$ admits a family of globally compatible product states?
@@
with base-change compatible incidence resolutions and a uniform singularity theory (Theorem 3.10). Moreover, the resulting intersection-theoretic and numerical data are fiberwise constant and computable (Theorem 5.5, Corollary 3.13). In this way, the background geometry controls which entanglement structures exist globally, and when they do exist, the family carries universal invariants that are stable under base change.
@@ 2.3.2. The Segre orbit inside the full Hilbert scheme of $\mathbb{P}^{n-1}$
(ii) Since $G_{\mathbf{d}}$ stabilizes $\Sigma_{\mathbf{d}}]$, the orbit map $\phi$ is constant on right cosets of $G_{\mathbf{d}}$, hence factors through a unique morphism $\bar{\phi}:\operatorname{PGL}_{n}/G_{\mathbf{d}}\to\operatorname{Hilb}(\mathbb{P}^{n-1})$.
@@ Appendix B Terms and notations
$\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\subset SB(\mathcal{A})$ When $r>1$, Schmidt-rank $\leq r$ locus. In general, $\Sigma_{\leq r}$ contains both entangled and pure states. When $r=1$, product-state locus.
@@
$\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})$ Relative incidence resolution of $\Sigma_{\leq k}(\mathcal{A},\mathbf{d})$.

### v2

@@ Abstract.
Quantum entanglement is a defining signature and resource of quantum theory, but its standard definition presupposes a globally fixed decomposition into subsystems. We develop a geometric framework that detects when such a decomposition cannot be globalized for twisted families of pure-state spaces. Using Severi–Brauer schemes associated to Azumaya algebras over a base scheme, we study pure-state entanglement in families of projective state spaces that are locally trivial but globally twisted. For a given factorization type, we show that the existence of a global locus of product states is equivalent to a reduction of the underlying projective linear torsor to the stabilizer of the corresponding Segre variety, so entanglement in families becomes a geometric obstruction to globalizing subsystem structure.
We construct the moduli space of subsystem structures, identify it with a natural torsor quotient, and realize it as a locally closed locus in the relative Hilbert scheme, yielding a canonical compactification by degenerations of product-state loci. In the bipartite case, once a subsystem structure is chosen, the Schmidt-rank stratification globalizes to a flat filtration with base-change compatible incidence resolutions and fiberwise constant numerical invariants. We complement this with Brauer-theoretic constraints and explicit examples showing that reducibility can depend on the underlying torsor rather than only on the Brauer class, with an interpretation via entangling monodromy.
@@ 1. Introduction
which is precisely the locus of product states of type $\mathbf{d}=(d_{1},\dots,d_{s})$. Entangled states are quantum states that do not belong to $\Sigma_{\mathbf{d}}$ and therefore exhibit non-classical correlations. By contrast, product states are uncorrelated. In quantum theory, entangled states are of particular interest.
@@
which is fppf-locally a trivial $\mathbb{P}^{n-1}$–bundle.
@@
    When does $SB(A)$ contain a relative Segre subscheme of type $\mathbf{d}$ (when does the associated $\mathrm{PGL}_{n}$-torsor admit a reduction to $G_{\mathbf{d}}$)? In terms of quantum information, when does $SB(\mathcal{A})$ admit a family of globally compatible product states?
@@
(0) The Severi–Brauer scheme as a twisted family of quantum states.
From the quantum-information viewpoint, $SB(\mathcal{A})\to X$ is a natural algebro–geometric model for a family of pure-state spaces whose projective trivializations may carry nontrivial twisting.
Indeed, even for a single $n$-level system the algebra of observables is the noncommutative algebra $\operatorname{End}(\mathcal{H})\cong\operatorname{Mat}_{n}$, whereas the space of pure states is the (commutative) projective space $\mathbb{P}(H)\cong\mathbb{P}^{n-1}$ of rays. The Severi–Brauer scheme globalizes this familiar dictionary over a parameter scheme $X$.
Since $\mathcal{A}$ is Azumaya, there exists an fppf cover $\{U_{i}\to X\}$ on which $\mathcal{A}$ splits:
$\mathcal{A}|_{U_{i}}\simeq\operatorname{End}(\mathcal{H}_{i})$
for some rank-$n$ vector bundle $\mathcal{H}_{i}$ on $U_{i}$. On each $U_{i}$, reduced rank-$1$ right ideals in $\operatorname{End}(\mathcal{\mathcal{}}H_{i})$ are parametrized by the projective bundle $\mathbb{P}(\mathcal{H}_{i})$: a line $l\subset(H_{i})_{x}$ corresponds to the right ideal $I_{l}:=\{\varphi\in\operatorname{End}(H_{i})\mid\operatorname{Im}(\varphi)\subset l\}$. Hence $SB(A)|_{U_{i}}\simeq\mathbb{P}(\mathcal{H}_{i})$ and the ray space of an $n$-dimensional quantum system.
The genuinely quantum feature is encoded not in the commutativity of $SB(\mathcal{A})$ but in the gluing. On overlaps $U_{ij}$ two splittings differ by an automorphism of $\operatorname{Mat}_{n}$, hence by a projective change of basis
$g_{ij}\in\mathrm{Aut}(\operatorname{Mat}_{n})\cong PGL_{n},$
which acts on rays and glues the local projective bundles. This means that the family of states can be realized as
$SB(\mathcal{A})\simeq P\times^{PGL_{n}}\mathbb{P}^{n-1}$
for the $PGL_{n}$-torsor $P\to X$ associated to $\mathcal{A}$. If the underlying $\operatorname{PGL}_{n}$-torsor $P\to X$ is nontrivial, the Brauer class $A]\in\operatorname{Br}(X)$ is nonzero. (Even when $[A]=0$, the projective bundle $SB(A)$ need not be trivial. It is then of the form $\mathbb{P}(E)$ for some rank-$n$ vector bundle $E$, unique up to tensoring by a line bundle.) Thus, while $SB(\mathcal{A})$ is a commutative scheme of pure states, the “quantumness” survives as twisting/monodromy encoded by $\mathcal{A}$. However, not all quantum states are of interest. Product states are valid quantum states, but they lack quantum correlations and can be regarded as classical. What is truly central to quantum theory are entangled states – those that cannot be expressed as product states. Our remaining objective is to study these states in the context of the Severi–Brauer scheme.
@@
with base-change compatible incidence resolutions and a uniform singularity theory (Theorem 3.10). Moreover, the resulting intersection-theoretic and numerical data are fiberwise constant and computable (Theorem 5.5, Corollary 3.13). In this way, the background geometry controls which entanglement structures exist globally, and when they do exist, the family carries invariants that are stable under base change.
@@ 2.3.2. The Segre orbit inside the full Hilbert scheme of $\mathbb{P}^{n-1}$
(ii) Since $G_{\mathbf{d}}$ stabilizes $\Sigma_{\mathbf{d}}]$, the orbit map $\phi$ is constant on left cosets of $G_{\mathbf{d}}$, hence factors through a unique morphism $\bar{\phi}:\operatorname{PGL}_{n}/G_{\mathbf{d}}\to\operatorname{Hilb}(\mathbb{P}^{n-1})$.
@@ Appendix B Terms and notations
$\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\subset SB(\mathcal{A})$ When $r>1$, Schmidt-rank $\leq r$ locus. In general, $\Sigma_{\leq r}$ contains both entangled and product states. When $r=1$, product-state locus.
@@
$\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})$ Relative incidence resolution of $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$.